A bag contains 4 gold marbles, 9 silver marbles, and 24 black marbles. The rules of the game are as follows: You randomly select one marble from the bag. If it is gold, you win $4, if it is silver, you win $3. If it costs $1 to play, what is your expected profit or loss if you play this game?
Answer #1

4 gold marbles, 9 silver marbles, and 24 black marbles

Total marbles = 4 + 9 + 24 = 37

P[ Gold marble ] = 4/37

Expected payoff of gold marble = $4 - $1 = $3 ( he will have to ay $1 to play )

P[ Silver marble ] = 9/37

Expected payoff of silver marble = $3 - $1 = $2 ( he will have to ay $1 to play )

P[ Black marble ] = 24/37

Expected payoff of black marble = $0 - $1 = - $1 ( he will have to ay $1 to play )

Expected profit or loss = P[ Gold marble ]*Expected payoff of gold marble + P[ Silver marble ]*Expected payoff of silver marble + P[ Black marble ]*Expected payoff of black marble

Expected profit or loss = (4/37)*3 + (9/37)*2 + (24/37)*(-1)

Expected profit or loss = 12/37 + 18/37 - 24/37

Expected profit or loss = 30/37 - 24/37

Expected profit or loss = 6/37

Expected profit or loss = $0.162

Exected Profit of $0.162
Answer #2

ANSWER :



Total marbles = 4 + 9 + 24 = 37 


My expected value

= P(gold) * 4 + P(silver) * 2  + P(black) * 0  - Cost of playing

= 4/37 * 4 + 9/37 * 3 + 0 - 1

= 6/37 = 0.1621  ($) (ANSWER)

answered by: Tulsiram Garg